# Refinement Module based on Parse Graph of Feature Map for Human Pose Estimation

### v1

@@ Abstract
Parse graphs of the human body can be obtained in the human brain to help humans complete the human pose estimation (HPE). It contains a hierarchical structure, like a tree structure, and context relations among nodes. Many researchers pre-design the parse graph of body structure, and then design framework for HPE. However, these frameworks are difficulty adapting when encountering situations that differ from the preset human structure. Different from them, we regard the feature map as a whole, similarly to human body, so the feature map can be optimized based on parse graphs and each node feature is learned implicitly instead of explicitly, which means it can flexibly respond to different human body structure. In this paper, we design the Refinement Module based on the Parse Graph of feature map (RMPG), which includes two stages: top-down decomposition and bottom-up combination. In the top-down decomposition stage, the feature map is decomposed into multiple sub-feature maps along the channel and their context relations are calculated to obtain their respective context information. In the bottom-up combination stage, the sub-feature maps and their context information are combined to obtain refined sub-feature maps, and then these refined sub-feature maps are concatenated to obtain the refined feature map. Additionally ,we design a top-down framework by using multiple RMPG modules for HPE, some of which are supervised to obtain context relations among body parts. Our framework achieves excellent results on the COCO keypoint detection, CrowdPose and MPII human pose datasets. More importantly, our experiments also demonstrate the effectiveness of RMPG on different methods, including SimpleBaselines, Hourglass, and ViTPose.
@@ I Introduction
Figure 1: (a) The parse graph of body structure. (b) The parse graph of feature maps. The red dotted lines in (a) and (b) are the context relations among sub-feature maps.
@@
When observing a person, humans can decompose the human body from the whole to part to primitives, thereby understanding the person more comprehensively. This structure can be expressed using parse graphs [1, 2]. As shown in Fig. 1a, the parse graph includes the hierarchical structure and context relations among structures [1]. The parse graph is usually infered through bottom-up and top-down ways [1, 3, 4], which is also consistent with human visual patterns [5, 6]. In order for the model to have such capabilities, researchers try to use parse graph in various visual tasks, such as human pose estimation [2, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], and achieve good results. These methods decompose the human body and then perform inference. However, since the features are fixed after decomposition, the framework lacks flexibility and can hardly accurately complete HPE when encountering different human body structures. Different from them, we treat the feature map as a whole, like the human body. Then according to the parse graph of the feature map, the feature map is optimized through top-down decomposition and bottom-up combination (see Fig. 1b). Since the features of nodes can be learned implicit, which means the features are not fixed, it can flexibly cope with the diversity and complexity of samples.
@@
In this paper, we design the Refinement Module based on the Parse Graph of feature map (RMPG) and use it to design a framework with a hierarchical structure, from body to parts to joints. For the RMPG, specifically, as shown in Fig. 1b, it consists of two stages, top-down decomposition and bottom-up combination. In the top-down decomposition stage, the feature map is decomposed along the channel dimension, and the context relations among sub-feature maps are calculated to obtain their respective contextual information. The sub-feature maps also can be segmented which is an expansion along the depth direction. In the bottom-up combination stage, the sub-feature maps containing context information are concatenated to obtain the optimized feature map. Since it is optimizing feature maps, it can be easily applied by other methods. In addition, the RMPG module can help explicitly obtain the context relations among body parts in Fig. 1a. It only supervises sub-feature maps without the need to design additional network modules or frameworks. This is demonstrated by our top-down framework, which consists of multiple RMPGs, some of which are supervised. In summary, the contributions of this paper are as follows:
@@
  • We propose a new Refinement Module based on the Parse Graph of feature map (RMPG). And the RMPG module effectiveness is demonstrated in other methods, including SimpleBaseline [21], Hourglass [22] and ViTPose [23].
  • We propose a top-down framework based on RMPG to explicitly modeling context relations and hierarchical structure in parse graph of body structure.
@@
  • Our framework is demonstrated on the CrowdPose, COCO keypoint detection, and MPII human pose datasets.
@@
Figure 2: (a) Our network framework follows a top-down prediction method, with high-level features guiding low-level feature learning, from body to limbs to joints. Through the RMPG module with supervision, the context relations among body parts can be obtained to help refine each body part (the yellow RMPG modules with supervision, and the blue RMPG modules without supervision). The heatmap in the red rectangular dotted box is not the direct feature map supervision information, but the feature map refined by the previous RMPG module to represent the refined structural features. The actual supervision information is located inside the RMPG module. (b) The RMPG module mainly operates on feature maps. First, the feature map is split into multiple sub-feature maps along the channel. Then calculate the context and self relations (inter-class and intra-class relations) of each child node to optimize each sub-node. The finally optimized child nodes are concatenated to obtain the refined feature map. (c) An example of extending the base module of RMPG along the depth direction.
Figure 3: (a) Supervision from body to limb. (b) Supervision from limb to joints.
@@ II Related work
Parse graphs. The parse graph includes a tree-structured decomposition in the vertical direction and context relations [1]. As shown in Fig. 1a, the human body is first divided into various parts, such as the head and left leg. In addition, each part can continue to be decomposed. Many methods use hierarchical structures similar to the parse graph to accomplish visual tasks, such as human pose estimation [15, 7, 8, 10, 11, 16, 17, 19, 20, 18], segmentation [24, 25, 26, 27], object detection [28, 24, 29] and human parsing [30]. These methods perform hierarchical parse of scenes or objects (e.g. human bodies) to better complete visual tasks. However, the actual situation is complex and there is no guarantee that the scene or object satisfies the proposed hierarchical structure, making it difficult for the model to effectively handle new samples outside the hierarchical structure., such as [2, 15, 7, 30, 29]. In addition, although some methods model hierarchical structures, they lack modeling of context relations, such as [24, 7, 28, 29, 26, 27]. Different from these methods, we treat the feature map as an object and then optimize the feature map based on the parse graph. As shown in Fig. 1b, the feature map is refined through top-down decomposition and bottom-up combination. Because it operates on feature maps, it can be easily used by other methods. It is also helpful in helping to obtain context relations in the parse graph of body structure. As shown in Fig. 2a, our framework uses multiple RMPGs to explicitly model hierarchical structure and context relations among different parts, with the latter obtained by supervising RMPG’s sub-feature maps. Experiments demonstrate the effectiveness of our framework and the generality of RMPG.
@@
CNN based HPE. Some methods [31, 22, 32, 33, 21, 34, 35, 36] design many excellent network architectures for HPE. This method of improving backbone feature extraction capabilities for HPE seems to have become mainstream. However, these methods only rely on appearance features for learning, which can easily lead to network overfitting [1]. Inspired by [2], our framework as shown in Fig. 2a, we use multiple RMPG modules to model context relations and hierarchy in the parse graph of body structure.
@@
Transformer based HPE. Transformer is proposed by [37] for neural machine translation (NMT) task [38]. It is widely used in various visual tasks and achieves good performance, such as HPE [23, 39, 40]. Most methods focus on the application of transformer [40, 23, 41] and simplifying transformer [42, 43]. On the contrary, our RMPG may provide possible optimization directions for the transformer. The transformer core is multi-headed attention and the heads are similar to nodes in a tree structure. However, compared with Fig. 1b, it lacks the context relations modeling for each head and does not further decomposition and combination for each head, which may be the direction of transformer [37] optimization. We add the RMPG module to the method of ViTPose [23]. Experimental results demonstrate the effectiveness of our module and also provide support for the possible optimization directions of the transformer structure mentioned above.
@@ III Method
In this section, the overall framework is described. Then the specific design of RMPG will be introduced in detail. Finally, we illustrate how we set up supervision in RMPG to obtain the context information among body parts. For the derivation of the parse graph of body structure, please see [2].
III-A Architecture
As shown in Fig. 2a, Our framework is modified based on HRNet [31], where $F_{1}$ is the largest scale feature map in the fourth stage. Our network framework use high-level features to guide low-level feature learning. We divide the human body into three levels, namely body, parts, and joints. And the context relations among different body parts is utilized to help refine predictions. In the Top-down process, we use the RMPG module with the supervision to obtain the context relations among different body parts and self relations. After obtaining the joint’s refined feature map $F_{3}$, we use the RMPG module without supervision to get the final prediction result.
@@ III-A Architecture
III-B The design of context and self relations in RMPG
The RMPG module includes top-down decomposition and bottom-up composition. As shown in Fig. 2b, The feature map $F\in R^{C\times H\times W}$ is sliced into $N$ sub-feature maps $F_{1},F_{2},...,F_{N}\in R^{\frac{C}{N}\times H\times W}$ along the channel dimension, where $N=3$. In order to calculate the context relations among sub-feature maps and self relations in each sub-feature map, We reshape the $N$ sub-feature maps into $R^{L\times\frac{C}{N}}$, $L\in H\times W$, respectively. Then we concatenate the reshaped sub-feature maps to get $F_{all}\in R^{LN\times\frac{C}{N}}$. Finally the context and self relation can be obtained:
$\begin{split}Rel=F_{all}\times F_{all}^{T}\end{split}$ (1)
where $\times$ represents matrix multiplication and $Rel\in R^{LN\times LN}$ is the spatial correlation of the sub-feature maps, including the self-correlation and inter-correlation. Then $F_{all}^{*}$ can be obtained:
$\begin{split}F_{all}^{*}=SoftMax(Rel)\times F_{all}\end{split}$ (2)
where $F_{all}^{*}\in R^{LN\times\frac{C}{N}}$ contains the context relations and self-attention relations of each sub-features. Then, we reshape $F_{all}^{*}$ in $R^{C\times H\times W}$. Finally, $F_{all}^{*}$ and $F$ are added to obtain the refined feature map. The above is the refinement method when the depth is 1. When the depth is greater than 1, such as 2, each sub-feature map needs to be optimized through decomposition and combination according to the above method. The optimized sub-features are then used to further refine the original input feature map. This process is similar to recursion.
III-C The design of supervision
We use the body heatmap to supervise the feature map $F_{1}$ (see Fig. 2a) and then the feature map is split and combined along the channel in RMPG (see Fig. 3a), and the sub-feature map after the first split is supervised using different body parts heatmaps, including left and right legs, left and right arms, and torso, a total of 5 parts, which are the same as [2]. Then, the refined feature maps $F_{2}$ (see Fig. 2a) of all body parts are refined again through the RMPG module, and different joint heatmaps are used to supervise the sub-feature maps after the first segmentation in RMPG (see Fig. 3b). The method and number of groups in the parse graph of body structure are the same as [2]. For the method of making heatmaps of different hierarchical structures, see [2]. Finally, the refined feature map $F_{3}$ (see Fig. 2a) is refined by the unsupervised RMPG module to obtain the final prediction result.
IV Experiments

### v3

@@ Abstract
Parse graphs of the human body can be obtained in the human brain to help humans complete the human Pose Estimation better (HPE). It contains a hierarchical structure, like a tree structure, and context relations among nodes. To equip models with such capabilities, many researchers predefine the parse graph of body structure to design HPE frameworks. However, these frameworks struggle to adapt to instances that deviate from the predefined parse graph and are often parameter-heavy. Unlike them, we view the feature map holistically, much like the human body. It can be optimized using parse graphs, where each node’s feature is an implicit expression rather than a fixed one. This allows it to adapt to more instances, unconstrained by rigid structural features. In this paper, we design the Refinement Module based on the Parse Graph of feature map (RMPG), which includes two stages: top-down decomposition and bottom-up combination. In the first stage, the feature map is decomposed into multiple sub-feature maps along the channel. In the second stage, the context relations of sub-feature maps are calculated to obtain their respective context information and the sub-feature maps with context information are concatenated along channels to obtain the refined feature map. Additionally, we design a hierarchical network with fewer parameters using multiple RMPG modules to model the context relations and hierarchies in the parse graph of body structure for HPE, some of which are supervised to obtain context relations among body parts. Our network achieves excellent results on multiple mainstream human pose datasets. More importantly, the effectiveness of RMPG is proven on different methods. The code of RMPG will be open.
@@ I Introduction
Figure 1: (a) The parse graph of body structure, from [1]. The human body is partitioned into 5 parts (limbs and torso) and structured into 3 hierarchical levels (body, parts, joints), with semantic granularity decreasing from top to bottom. (b) The parse graph of feature maps.
@@
When observing a person, humans can decompose the human body from the whole to part to primitives, thereby understanding the person more comprehensively. This structure can be expressed using parse graphs [2, 1]. As shown in Fig 1a, the parse graph includes the hierarchical structure and context relations among structures [2]. Hierarchical structure, as a tree-like decomposition, helps models capture multi-level relations, while context relations describe spatial interactions between parts to ensure spatial consistency. The parse graph is usually infered through top-down and bottom-up ways [2, 3, 4], which is also consistent with human visual patterns [5, 6]. In order for the model to have such capabilities, many methods [7, 8, 9, 10, 11, 12, 13, 14, 15, 16] try to use the context information or hierarchy in the parse graph of body structure for HPE. Liu et al. [1] think that the above methods do not model context relations and hierarchical structures simultaneously. So, they designed a new network to model context relations and hierarchies in the parse graph of body structure for HPE and achieve good results. However, its fixed body parse graph struggles with diverse human poses, and its more parameters limit its applicability. Different from them, we treat the feature map as a whole, like the human body. Then according to the parse graph of the feature map, the feature map is optimized through top-down decomposition and bottom-up combination (see Fig. 1b). Since the features of nodes can be learned implicit, which means the features are not fixed, it mitigates, to some extent, the problem of poor adaptability to various human poses.
@@
In this paper, we design the Refinement Module based on the Parse Graph of feature map (RMPG) and use it to design a hierarchical network with fewer parameters than [1] to model the context relations and hierarchical in the parse graph of body structure (see Fig. 1a). For the RMPG, specifically, as shown in Fig. 1b, it consists of two stages, top-down decomposition and bottom-up combination. In the top-down decomposition stage, the feature map is decomposed along the channel dimension. The sub-feature maps also can be segmented which is an expansion along the depth direction. In the bottom-up combination stage, the context relations of sub-feature maps are calculated to obtain their respective context information and the sub-feature maps containing context information are concatenated to obtain the optimized feature map. Since it is optimizing feature maps, it can be easily applied by other methods. In addition, the RMPG module can help explicitly model the context relations and hierarchies in Fig. 1a. It only supervises sub-feature maps without the need to design additional network modules or frameworks. This is demonstrated by our hierarchical network, which consists of multiple RMPGs, some of which are supervised. In summary, the contributions of this paper are as follows:
@@
  • We propose a new Refinement Module based on the Parse Graph of feature map (RMPG) to partially solve the problem that fixed body parse graphs are difficult to adapt to different samples. And the RMPG module effectiveness is demonstrated in other methods, including SimpleBaseline [17], Hourglass [18] and ViTPose [19].
  • We propose a hierarchical network based on RMPG to address the issue of excessive parameters when explicitly modeling contextual relationships and hierarchical structures in parse graphs.
@@
  • Our hierarchical network is demonstrated on the CrowdPose, COCO keypoint detection, and MPII human pose datasets.
@@
Figure 2: (a) Our network uses high level features to guide low level feature learning from the body to parts to joints, where $F_{1}$ comes from HRNet [20]. Through the RMPG module with supervision, the context relations among body parts can be obtained to help refine each body part (the yellow RMPG modules with supervision, and the blue RMPG modules without supervision). The heatmap within the red dashed box represents refined structural features from the previous RMPG module, with actual supervision information residing inside the RMPG module. Note: All heatmap supervision on feature maps is performed after dimensionality processing through a conv2d layer. (b) The RMPG module mainly operates on feature maps. First, the feature map is split into multiple sub-feature maps along the channel. Then the context relation of each child node is calculated to optimize each sub-node. The finally optimized child nodes are concatenated to obtain the refined feature map.
Figure 3: (a) Supervision from body to parts. (b) Supervision from parts to joints.
@@ II Related work
Parse graphs. The parse graph includes a tree-structured decomposition in the vertical direction and context relations [2]. As shown in Fig. 1a, the human body is first divided into five parts, including limbs and torso. Furthermore, each part can continue to be decomposed. Many methods use hierarchical structures and context information in object parse graphs or other structure restraint to accomplish visual tasks, such as human pose estimation [21, 22, 23, 1, 15, 7, 8, 10, 11, 16], segmentation [24, 25, 26] and object detection [27, 24, 28]. Some of these methods [23, 24, 7, 27, 28, 25, 26, 15, 8, 10, 11, 16, 27] fail to simultaneously model the context relations and hierarchical structures in object parse graphs. Although PGBS [1] solves this problem with high parameters, it relies on fixed human structure splits and combinations, making it difficult to adapt to diverse postures. In addition, Chen et al. [22] use a discriminator to distinguish true and false postures, but training is difficult and Li et al. [21] use Transformer [29] to learn position relationship constraints, but its performance is limited on the MPII dataset [30], which may be because more data (e.g., the COCO keypoint detection dataset [31]) is required to achieve better results after introducing the Transformer. Different from these methods, we treat the feature map as an object and then optimize the feature map based on the parse graph. As shown in Fig. 1b, because it operates on feature maps, it can be easily used by other methods. In addition, as shown in Fig. 2a, RMPG also helps our network model context relations and hierarchies in the parse graph of body structure with lower parameters. Experiments demonstrate the effectiveness of our network and the generality of RMPG.
@@
CNN based HPE. Some methods [20, 18, 32, 33, 17, 34, 35, 36] design many excellent network architectures for HPE. This method of improving backbone feature extraction capabilities for HPE seems to have become mainstream. However, these methods only rely on appearance features for learning, which can easily lead to network overfitting due to lacking the information of structure [2]. Inspired by [1], our hierarchical network as shown in Fig. 2a where $F_{1}$ comes from HRNet [20], we use multiple RMPG modules to model context relations and hierarchy in the parse graph of body structure.
@@
Transformer based HPE. Transformer is proposed by [29] for neural machine translation (NMT) task [37]. It is widely used in various visual tasks and achieves good performance, such as HPE [19, 38, 21, 39]. Most methods focus on the application of transformer [39, 19, 40, 21] and simplifying transformer [41, 42]. On the contrary, our RMPG may provide possible optimization directions for the transformer. The transformer core is multi-headed attention and the heads are similar to nodes in a tree structure. However, compared with Fig. 1b, it lacks the context relations modeling for each head and does not further decomposition and combination for each head, which may be the direction of transformer [29] optimization. We add the RMPG module to the method of ViTPose [19]. Experimental results demonstrate the effectiveness of our module and also provide support for the possible optimization directions of the transformer structure mentioned above.
@@ III Method
In this section, we first describe the derivation of the parse graph. Then, we present the overall network architecture. Subsequently, we detail the specific design of RMPG. Finally, we explain how to set up supervision in RMPG to capture context information between body parts.
III-A Parse graph
The parse graph includes hierarchical structures and contextual relations. It is represented as a 4-tuple $(V,E,\psi^{and},\psi^{leaf})$, where $(V,E)$ defines the graph structure, and $(\psi^{and},\psi^{leaf})$ are potential functions. Each node $u\in V$ has state variable $s_{u}=\{x_{u},y_{u}\}$, where $x_{u}$ is the position and $y_{u}$ is the type. The probability of the state variables $\Omega$ given an image $I$ is:
$P(\Omega|I)=\frac{1}{Z}\exp\{-E(\Omega,I)\}$ (1)
where $E(\Omega,I)$ is the energy function and $Z$ is the partition function. The energy function $F(\Omega)=-E(\Omega,I)$ is decomposed as:
$F(\Omega)=\sum_{u\in V_{L}}\psi^{leaf}_{u}(s_{u},I)+\sum_{u\in V_{A}}\psi^{and}_{u}(s_{u},{s_{v}}_{v\in C(u)})$ (2)
where $V_{L}$ and $V_{A}$ are leaf and non-leaf nodes respectively, and $C_{u}$ denotes the children of node $u$. The optimal state $\Omega^{*}$ is computed in two stages: bottom-up activation and top-down refinement. The bottom-up stage computes the maximum score $F_{u}^{\uparrow}(s_{u})$, while the top-down stage refines each node $v$ using its parent node $u$ and siblings:
$F_{v}^{\downarrow}(s_{v})=\psi_{u,v}(s_{u}^{*},s_{v})+\xi_{v}(s_{v},{s_{h}}_{h\in S_{v}})$ (3)
Where $S_{v}$ contains all nodes at the same level as $v$, $\xi_{v}$ captures context relations, and $s_{u}^{*}=\arg\max_{s_{u}}F_{u}^{\uparrow}(s_{u})$. This two-stage process ensures accurate part predictions by leveraging hierarchical and context information.
III-B Architecture
Our network explicitly models the context relations and hierarchical structure in Fig. 1a. As shown in Fig. 2a, our network is modified based on HRNet [20], where $F_{1}$ is the largest scale feature map in the fourth stage of HRNet. $F_{1}$ is used to generate the human body heatmap after a 2D convolution, which is supervised during training. We divide the human body into three levels, namely body, parts, and joints. For the parts and joints, we use the RMPG module with the supervision to obtain the context relations (see Fig. 3). After obtaining the refined feature map $F_{3}$ of the joint, we then use the unsupervised RMPG module to refine $F_{3}$ to obtain $F_{4}$. Finally, $F_{4}$ is used to generate the final joint heatmaps after a 2D convolution, which is supervised during training.
@@ III-B Architecture
Figure 4: (a) An input image of our network. (b) Visualization of context relations among all joint in (a). Each row represents the relation between a corresponding joint and all joints. The higher the score, the stronger the relation. From the visualization of the context relations in joints we notice that the joint tends to have a higher score with itself when it is visible , such as left hip, left wrist and left ankle. On the contrary, the joints tends to have a higher score with other joints when it is occluded or invisible, such as right wrist, right ankle and right shoulder.
III-C The design of RMPG
The RMPG module includes top-down decomposition and bottom-up composition, propagating context information from leaf nodes to higher levels to optimize feature maps, inheriting the hierarchical structure and context relations in Sec 3.1. The decomposition rule is to split the parent node features into individual child node features along the channel, which is a recursive operation. The combination rule is to recursively concatenate the child nodes with context information along the channel, which is a recursive operation.
@@ III-C The design of RMPG
As shown in Fig. 2b, the feature map $F\in R^{C\times H\times W}$ is sliced into $N$ sub-feature maps $F_{1},F_{2},...,F_{N}\in R^{\frac{C}{n}\times H\times W}$ along the channel dimension, where $n=N=3$. Note that $n$ and $N$ can also be different (e.g., Eq. 8). In order to calculate the context relations among sub-feature maps in each sub-feature map, We reshape the $N$ sub-feature maps into $R^{L\times\frac{C}{n}}$, $L\in H\times W$, respectively. Then we concatenate the reshaped sub-feature maps to get $F_{all}\in R^{LN\times\frac{C}{n}}$. Finally the Context Information $CI()$ can be obtained:
$CI(\{F_{1},F_{2},\cdots,F_{N}\})=SoftMax(E)\times F_{all}$ (4)
where $E=F_{all}\times F_{all}^{T},E\in R^{LN\times LN}$ is the spatial correlation of the sub-feature maps, including the intra-correlation and inter-correlation and $\times$ represents matrix multiplication. $CI(\{F_{1},F_{2},\cdots,F_{N}\})\in R^{LN\times\frac{C}{n}}$ contains the context relations relations of each sub-features. Then, $CI(\{F_{1},F_{2},\cdots,F_{N}\})$ is reshaped in $R^{C\times H\times W}$. The refinement of the above RMPG module is completed. Finally, to preserve the properties of the original feature $F$, we add it as a residual to $CI$. The $CI()$ is $\xi_{v}=(s_{v},{s_{h}}_{h\in S_{v}})$ in Eq. 3.
The above is the refinement method when the depth is 1. When the depth is greater than 1, such as 2 (see Fig. 2c), the optimization process is as follows. First, there is the top-down decomposition process. The input feature map $F\in R^{C\times H\times W}$ will be decomposed to obtain all node features according to the setting of $gp=[3,3]$ (see Sec. 4.2):
$\begin{split}N=\{F,V_{1}^{1},V_{1}^{2},V_{1}^{3},V_{0}^{1},V_{0}^{2},\cdots\}\end{split}$ (5)
where $N$ is the set of features of all nodes in the parse graph of $F$:
$ch(V_{1}^{j})$ =$=\{V_{0}^{3(j-1)+k}\}_{k=1}^{3},\quad j=1,2,3$ (6)
$ch(F)$ =$=\{V_{1}^{1},V_{1}^{2},V_{1}^{3}\}$ (7)
where $ch$ represents child nodes. Child nodes are obtained by decomposing the parent node along the channel, so $V_{0}^{i}\in R^{\frac{C}{9}\times H\times W},i={1,2,\cdots,9}$ and $V_{1}^{m}\in R^{\frac{C}{3}\times H\times W},m={1,2,3}$. Then there is the bottom-up combination process. The context relations between leaf nodes belonging to the same parent node are calculated according to Eq. 4:
$C_{0}^{j}=CI(ch(V_{1}^{j})),\quad j=1,2,3$ (8)
where $C_{0}^{j}\in R^{L3\times\frac{C}{9}}$ is reshaped in $R^{\frac{C}{3}\times H\times W}$, which can be viewed as the node $V_{1}^{j}$ with context information, represented by $V_{1}^{j^{*}}$. Then the context information among $V_{1}^{j^{*}}$ is also calculated:
$F_{r}=CI(\{V_{1}^{j^{*}}\}_{j=1}^{3})$ (9)
where $F_{r}\in R^{L3\times\frac{C}{3}}$ is reshaped in $R^{C\times H\times W}$. The refinement of the above RMPG module is completed. Finally, to preserve the properties of the original feature $F$, we add it as a residual to $F_{r}$. The reasoning for other different $gp$ settings is similar to the above.
@@
III-D Supervision
Supervision in our network. Firstly, we use the body heatmap to supervise the feature map $F_{1}$ after a 2D convolution (see Fig. 2a). Then, $F_{1}$ is decomposed (see Fig. 3a) according to the setting of $gp=[5,2]$:
$\begin{split}N=\{V_{n},V_{n-1}^{1},V_{n-1}^{2},\cdots\}\end{split}$ (10)
where $N$ is the set of nodes in the parse graph of $F_{1}$ and $n$ is the depth of the parse graph $n=len(gp)=2$. Next, perform supervision on all child nodes on the level 1:
$\begin{split}SV=\{V_{1}^{1},V_{1}^{2},V_{1}^{3},V_{1}^{4},V_{1}^{5}\}\end{split}$ (11)
where $V_{1}^{1}$ and $V_{1}^{2}$ are supervised by the left and right legs respectively, $V_{1}^{3}$ and $V_{1}^{4}$ are supervised by the left and right limbs respectively, and $V_{1}^{5}$ is supervised by torso. In this way, the context relations between parts can be obtained. Then, the refined feature maps $F_{2}$ (see Fig. 2a) are obtained, which is the result of the refinement containing all parts. Perform the same operation on $F_{2}$ as $F_{1}$ (see Fig. 3b), the difference is that the supervision information becomes joints. For example, $V_{1}^{1}$ is supervised by all joints on the left leg, including the left knee, left ankle, and left hip, and $V_{1}^{3}$ is supervised by all joints on the left limb, including the left shoulder, left elbow, and left wrist. Then, the refined feature maps $F_{3}$ (see Fig. 2a) are obtained, which is the result of the refinement containing all joints. Finally, $F_{3}$ is refined by the unsupervised RMPG module to obtain the final refinement feature maps $F_{4}$. Then $F_{4}$ goes through a 2d convolution to get the final joints prediction result.
Designing supervision. Follow the method of PGBS [1], the body heatmap is generated by placing a Gaussian kernel centered at the ground truth bounding box of the body and the size of the Gaussian kernel is proportional to the size of the human body in the image, and the parts heatmaps are generated by placing Gaussian kernels at the midpoints of skeletal segments, with kernel sizes proportional to bone lengths. For example, the left leg heatmap includes Gaussian kernels at the midpoints of the left hip-left knee and left knee-left ankle segments. During training, the score graphs $F_{v}^{\downarrow}(s_{v})$ in Eq. 3, which represent the heatmaps, are used as the ground truth to supervise our network.
IV Experiments